# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v1

@@ Appendix B $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [vempala2005geometric, Section 5].
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$. Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\|\pi_{t}-\pi\|_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$. If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ only needs $\mathcal{O}(d^{2}\log d\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon})$ iterations.
@@
[kannan1997random] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$. Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [kannan1997random, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, $\mathsf{Ball\ walk}$ often refers to ‘$\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of $\mathsf{Ball\ walk}$.
@@
$\lambda:=\mathbb{E}_{\pi}[\ell]=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x\,.$
Proposition 1 ([kannan1997random]).
@@
Proposition 2 ([kannan1997random, Page 22]).
$\|\nu-\pi\|_{\mathsf{TV}}\leq\frac{1-\lambda}{\lambda}$.
@@
Proposition 3 (Bound on average conductance, [kannan1997random, Corollary 4.5]).
@@
The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [kannan2006blocking] devising the blocking conductance and using the log-Cheeger inequality. When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [lee2017eldan] improves the mixing bound via the log-Cheeger constant.
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$. Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\|\nu_{t}-\nu\|_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$. From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{\mathcal{O}}(tM)$. When $\nu$ is isotropic, $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
Then, [kannan1997random] uses the following post-processing step to obtain an approximately uniform distribution on $\mathcal{K}$, with a provable guarantee.
@@
Proposition 4 ([kannan1997random, Theorem 4.16]).
Under the same setting above, assume $\|\nu_{t}-\nu\|_{\mathsf{TV}}\leq\varepsilon$ for step size $\delta\leq(8d\log\frac{d}{\varepsilon})^{-1/2}$ and fixed $t\in\mathbb{N}$. For $\bar{\nu}=\mathsf{law}(\bar{X})$ given by $\mathcal{A}$, it holds that $\|\bar{\nu}-\pi\|_{\mathsf{TV}}\leq\varepsilon$, and the expected number of calls on the conversion algorithm is at most $2$.
@@ Appendix C Functional inequalities
We provide full details on functional inequalities omitted in Section 3.1. We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and log-concave probability measure over $\mathbb{R}^{d}$, respectively.
@@
[cheeger1970lower] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [milman2009role], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [ledoux1994simple]., and then [kannan1995isoperimetric] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\|X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\|^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}(\Sigma_{\mu_{\mathsf{LC}}}))^{1/2}}\,.$ (C.1)
This immediately leads to $C_{\mathsf{PI}}(\pi)\lesssim(\mathbb{E}_{\pi}[\|X-\mathbb{E}_{\pi}X\|^{2}])^{1/2}\leq D^{2}$ for the uniform distribution $\pi$ over a convex body $\mathcal{K}$ with diameter $D>0$.
Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\|\cdot\|_{2}$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\|\Sigma_{\mu_{\mathsf{LS}}}\|_{2}^{1/2}}\,.$
While the original result in [kannan1995isoperimetric] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case. Following a line of work [lee2017eldan, chen2021almost, klartag2022bourgain, klartag2023logarithmic], the current bound is
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\|\Sigma_{\mu_{\mathsf{LS}}}\|_{2}^{1/2}}\,,$
@@
[ledoux1994simple] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ [^6] The opposite direction holds under dimension-scaling due to [ledoux1994simple]: $C_{\mathsf{LSI}}(\mu)\gtrsim C_{\mathsf{logCh}}^{-2}(\mu)/d$., and [kannan2006blocking] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$. Later in 2016, [lee2017eldan] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy. Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
@@ Appendix D The Wasserstein geometry
We present additional technical background on the Wasserstein geometry and Markov semigroup theory. Interested readers can refer to [villani2009optimal, ambrosio2005gradient, chewi2023log] for standard references on Wasserstein spaces and applications to sampling.
@@
While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot|z)$ if one starts (D.3) at $X_{0}=z$. This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform. The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [klartag2021spectral] as a reference to its application in this context.

### v2

@@ Appendix B $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [Vem05, §5].
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$. Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\lVert\pi_{t}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$. If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ needs $\mathcal{O}(d^{2}\frac{M^{2}}{\varepsilon^{2}}\log d\log\frac{M}{\varepsilon})$ iterations.
@@
[KLS97] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$. Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, $\mathsf{Ball\ walk}$ often refers to ‘$\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of $\mathsf{Ball\ walk}$.
@@
$\lambda:=\mathbb{E}_{\pi}\ell=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x\,.$
Proposition 1 ([KLS97]).
@@
Proposition 2 ([KLS97, Page 22]).
$\lVert\nu-\pi\rVert_{\mathsf{TV}}\leq\frac{1-\lambda}{\lambda}$.
@@
Proposition 3 (Bound on average conductance, [KLS97, Corollary 4.5]).
@@
The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [KLM06] devising the blocking conductance and using the log-Cheeger inequality. When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [LV17] improves the mixing bound via the log-Cheeger constant.
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$. Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$. From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{\mathcal{O}}(tM)$. When $\nu$ is isotropic, $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
Then, [KLS97] uses the following post-processing step to obtain an approximately uniform distribution on $\mathcal{K}$, with a provable guarantee.
@@
Proposition 4 ([KLS97, Theorem 4.16]).
Under the same setting above, assume $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for step size $\delta\leq(8d\log\frac{d}{\varepsilon})^{-1/2}$ and fixed $t\in\mathbb{N}$. For $\bar{\nu}=\mathsf{law}(\bar{X})$ given by $\mathcal{A}$, it holds that $\lVert\bar{\nu}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$, and the expected number of calls on the conversion algorithm is at most $2$.
@@ Appendix C Functional inequalities
We provide full details on functional inequalities omitted in §3.1. We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and log-concave probability measure over $\mathbb{R}^{d}$, respectively.
@@
[Che70] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [Mil09], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [Led94]., and then [KLS95] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\lVert X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\rVert^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}\Sigma_{\mu_{\mathsf{LC}}})^{1/2}}\,.$ (C.1)
This immediately leads to $C_{\mathsf{PI}}(\pi)\lesssim(\mathbb{E}_{\pi}[\lVert X-\mathbb{E}_{\pi}X\rVert^{2}])^{1/2}\leq D^{2}$ for the uniform distribution $\pi$ over a convex body $\mathcal{K}$ with diameter $D>0$.
Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\lVert\cdot\rVert_{2}$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,.$
While the original result in [KLS95] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case. Following a line of work [LV17, Che21, KL22, Kla23], the current bound is
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,,$
@@
[Led94] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ and $C_{\mathsf{LSI}}(\mu_{\mathsf{LS}})\gtrsim C_{\mathsf{logCh}}^{-2}(\mu_{\mathsf{LS}})$. [KLM06] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$. Later in 2016, [LV17] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy. Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
@@ Appendix D The Wasserstein geometry
We present additional technical background on the Wasserstein geometry and Markov semigroup theory. Interested readers can refer to [Vil09, AGS05, Che24] for standard references on Wasserstein spaces and applications to sampling.
@@
While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot|z)$ if one starts (D.3) at $X_{0}=z$. This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform. The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [KP21] as a reference to its application in this context.